# Scaling Low-Resource MT via Synthetic Data Generation with LLMs

### v1

@@ 5.3 Comparison with HPLT v2
To further assess our dataset’s utility, we conduct comparative experiments against HPLT v2 Burchell et al. (2025), a "real" parallel corpus derived from web sources (Internet Archive [^10] https://archive.org/ and Common Crawl  [^11] https://commoncrawl.org/).
We train systems for three out of the four overlapping language pairs (English paired with Basque, Icelandic and Macedonian). This decision was motivated by the significant variation in dataset sizes (Table 8), with Ukrainian having approximately ten times more data than the others, therefore, we leave it out. We train three models: (1) the same synthetic baseline as described earlier (synthetic), (2) a model trained on HPLT dataset (HPLT), and (3) a model trained on the concatenation of our synthetic dataset and the HPLT. All models are evaluated using ChrF on the same test set described previously. Table 4 reports the detailed ChrF scores.
@@
The models trained on our synthetic data alone perform on the same ballpark as the ones trained on the HPLT dataset, with an average difference of 2.23 ChrF points. The largest performance difference is observed for Macedonian, following a similar pattern as our previous experiments. These results demonstrate that our synthetic dataset is of sufficiently high quality to rival real-world parallel corpora, even when trained from scratch.
@@ 5.4 Beyond English-Centric MT: The Finnish Use Case
To explore the multilingual potential of our expanded dataset via pivoting (Section 3.4) and move beyond English-centric translation, we train MT models for two additional language pairs: Finnish–Somali and Finnish-Ukrainian. These languages were selected due to their prominence in Finland’s linguistic landscape, where Ukrainian and Somali are among the most widely spoken foreign languages, accounting for approximately 0.7% and 0.5% of the population, respectively Official Statistics of Finland (2024) (OSF). Developing high-quality models for these pairs is therefore both practical and relevant.
@@
NLLB-ft 42.26 42.31 47.61 51.62
[…]
Table 5: ChrF scores for Finnish-Somali and Finnish-Ukrainian translation. For each system, we report the base and fine-tuned (ft) variants, and the improvement ($\Delta$) from fine-tuning. “*” indicates a significant difference ($p<0.05$) under paired t-test and bootstrap resampling (5,000 iterations).
@@ 6 SynOPUS: a New Synthetic Parallel Corpus Repository
The increasing adoption of LLMs in generating synthetic data underscores the growing need to systematically organize synthetic datasets. Although it is well known that many parallel datasets already contain MT content Thompson et al. (2024), when synthetic data is intentionally produced, especially when involving significant financial or computational resources, proper archiving are paramount for promoting reuse, ensuring transparency, and maximizing resource utility. Therefore, with the release of our dataset, we introduce SynOPUS [^12] https://github.com/Helsinki-NLP/synOPUS, a new repository for parallel synthetic datasets, i.e., data that has been (partially) generated by translating text into other languages using MT systems or LLMs. We invite the community to contribute with their own datasets.
@@ 7 Conclusions
In this work, we thoroughly studied the quality and usefulness of LLM-generated synthetic data for low-resource MT. We presented a new synthetic corpus at document-level by forward translating Europarl, a parliamentary corpus, with GPT-4o. Then, we evaluated the resulting dataset both quantitatively and through human evaluation. Furthermore, we investigated the usefulness of this dataset for low-resource MT by: (i) identifying the most effective strategy for training, (ii) comparing our dataset with the public HPLT dataset, and (iii) extending our analysis beyond English-centric MT by generating a multi-parallel corpus via pivoting through English.
@@
LLMs are known to generate hallucinated content, outputs that are fluent and well-formed but factually incorrect or irrelevant Vázquez et al. (2025). This phenomenon is a risk in iself, as it can introduce noise and propagate misinformation in downstream MT models. In our case, we observed that in some cases, the model disregards the input and instead generates a response similar to, “You have been trained on data up until October 2023” in the target language. This issue is most prevalent in Georgian, with around 9,000 cases, and Ukrainian, with approximately 4,000 cases. For these languages, we removed each line containing the string “2023”. While such hallucinations appear to be an intrinsic limitation of current LLMs, they highlight the need for careful post-processing and validation when using synthetic data.
Data Contamination
Due to the the closed-source nature of GPT-4o, there is a risk of data contamination, since the model may have already seen our test set (FLORES-200) during pretraining. Recent studies Mansurov et al. (2024) have shown that distilled data may inherit intrinsic biases from the teacher model and this may have an impact on benchmark results. While this risk is inherent to most widely used test sets and cannot be fully controlled, we acknowledge it here for transparency.
@@ Appendix A Pilot Evaluation Results
OPUS-MT GPT EMMA Language Code Sent. Sent. Chunk Sent./3 Sent./0 Tosk Albanian als_Latn 53,9 61,3 63,1 44,7 54,4 Aragonese arg_Latn - 43,9 47,5 38,1 39,8 Aranese arn_Latn - 41,2 45,4 16,7 27,6 Asturian ast_Latn 59,6 59,6 62,8 52,7 53,3 Bashkir bak_Cyrl 40,6 51,4 52,2 24,2 29,1 Belarusian bel_Cyrl 44,4 46,6 49,2 39,0 37,9 Bosnian bos_Latn 58,8 63,9 65,4 51,6 53,3 Catalan cat_Latn 67,4 67,9 69,8 56,7 56,5 Crimean Tatar crh_Latn 35,9 37,6 40,0 13,7 23,2 Welsh cym_Latn 64,9 73,3 73,6 59,4 59,8 Esperanto epo_Latn 59,7 63,2 65,6 60,8 60,5 Basque eus_Latn 55,0 57,3 60,4 41,5 44,0 Friulian fur_Latn 49,9 45,7 50,0 45,4 43,8 Scottish Gaelic gla_Latn 42,6 52,4 55,7 44,5 45,9 Irish gle_Latn 60,5 61,3 61,6 48,9 49,3 Galician glg_Latn 62,5 63,3 66,2 55,7 55,9 Hebrew heb_Hebr 61,3 58,1 59,2 41,1 47,0 Croatian hrv_Latn 61,5 60,9 62,2 50,9 51,2 Armenian hye_Armn 47,5 56,3 56,5 45,7 48,7 Icelandic isl_Latn 53,0 55,6 59,3 41,1 41,6 Georgian kat_Geor 42,7 51,3 42,4 45,3 45,9 Ligurian lij_Latn 43,9 36,2 40,1 34,5 35,3 Limburgish lim_Latn 36,5 43,7 44,6 28,3 29,8 Lombard lmo_Latn 34,1 35,3 38,8 25,3 28,4 Luxembourgish ltz_Latn 55,5 59,4 60,9 51,5 49,9 Macedonian mkd_Cyrl 64,5 65,3 64,2 55,2 56,8 Occitan oci_Latn 64,0 67,5 67,0 51,7 46,0 Sicilian scn_Latn 39,9 45,0 48,7 43,4 43,8 Somali som_Latn 44,2 47,6 53,0 42,7 38,6 Sardinian srd_Latn 50,1 43,4 46,2 40,0 48,8 Serbian srp_Cyrl 63,0 64,1 65,0 35,5 46,7 Tatar tat_Cyrl 43,0 53,6 54,9 21,1 33,9 Turkmen tuk_Latn 42,6 55,1 54,8 22,3 25,8 Turkish tur_Latn 62,8 66,1 67,1 36,8 34,5 Uyghur uig_Arab 37,1 38,1 36,8 30,2 31,1 Ukrainian ukr_Cyrl 60,1 60,1 62,1 48,7 49,4 Northern Uzbek uzn_Latn 12,7 59,3 60,7 30,5 44,4 Venetian vec_Latn 44,6 49,5 52,0 34,7 37,1 Yiddish ydd_Hebr 0,0 40,5 42,6 41,4 51,5 Average 49,2 53,9 55,6 40,8 43,6
Table 6: ChrF scores of the evaluation of GPT and EMMA on low-resource MT. We compare performance at the sentence (Sent.) and chunk (Chunk) level. For EMMA, both zero-shot (Sent./0) and three-shot (Sent./3) settings are reported. The best OPUS-MT model is included as a reference. Highlighted rows correspond to the final set of selected languages for our study.
@@ Appendix C Training Regimes
  • Baseline: We employ a shared 32k SentencePiece Kudo and Richardson (2018) vocabulary trained on the synthetic corpus; other settings follow the original Transformer-base recipe. Mini-batch fitting is enabled to optimize memory usage. Validation every 2500 updates checks perplexity. Early-stopping is employed on the development set, with a patience of 10.
@@
  • NLLB-200-distilled-1.3B-ft: We fine-tune the NLLB-distilled-1.3B model with DeepSpeed on four V100 GPUs in FP16 mixed precision. Training uses a per-GPU batch size of 32 sentences, a maximum sequence length of 128 tokens, the Adam optimiser with a 1 ×$\times 10^{-4}$ learning rate, and runs for up to four epochs. DeepSpeed ZeRO-1 is used for basic tensor sharding; everything else is left on-GPU. Early-stopping is employed on the development set, with a patience of 5.
@@ Appendix D OPUS-MT Models selected for fine-tuning
We select the best available OPUS-MT model based on the OPUS-MT Dashboard Tiedemann and De Gibert (2023), by looking at the BLEU score on the FLORES200 dataset.
@@ Appendix E COMET scores for MT training
Model Language Pair ##\## Params
@@
[…]
Table 7: COMET scores of Section 5.1.
Appendix F Detailed Data Sizes and ChrF Scores for HPLT Comparison

### v2

@@ 5.3 Comparison with HPLT v2
To further assess our dataset’s utility, we conduct comparative experiments against HPLT v2 de Gibert et al. (2024); Burchell et al. (2025), a "real" parallel corpus derived from web sources (Internet Archive [^15] https://archive.org/ and Common Crawl). [^16] https://commoncrawl.org/
We train systems for three out of the four overlapping language pairs (English paired with Basque, Icelandic and Macedonian). This decision was motivated by the significant variation in HPLT v2 data sizes (see Appendix F, Table 10), with Ukrainian having approximately ten times more data than the others; therefore, we leave it out. We train three models: (1) the same synthetic baseline as described earlier (Synthetic), (2) a model trained on HPLT dataset (HPLT), and (3) a model trained on the concatenation of our synthetic dataset and the HPLT (HPLT + Synthetic). All models follow the same architecture (transformer-base) and hyperparameters. All models are evaluated using ChrF on the same test set described previously. Table 4 reports the detailed ChrF scores.
@@
The models trained on our synthetic data alone perform on the same ballpark as the ones trained on the HPLT dataset, with an average difference of 2.23 ChrF points. The largest performance difference is observed for Macedonian, following a similar pattern as our experiments in the previous section. These results demonstrate that our synthetic dataset is of sufficiently high quality to challenge real-world parallel corpora, even when trained from scratch.
@@
5.4 Effect of Training Data Size
To assess the efficiency and scalability of our synthetic data approach, we train models with increasing fractions of the available synthetic data (10–100%) by creating cumulative subsets. For each subset, we trained a Marian NMT model using the same tokenizer and identical hyperparameters to the baselines presented in Section 5.1 to ensure comparability across runs, where data is the only changing variable.
Effective Scaling with Synthetic Data
Figure 2 shows that performance improves consistently as more data is used, confirming that additional synthetic data is beneficial across all language directions. However, the gains decrease after 50–60% has been used. For example, English-Basque improves by +8.7 chrF from 10% to 50%, but only +1.4 additional points from 50% to 100%. This means that larger synthetic corpora brings gains but substantial improvements can already be achieved with a fraction of the full dataset. This suggests that future research on synthetic data generation should prioritize data quality and diversity for greater benefits than further scaling alone.
Figure 2: Learning curves with different data sizes.
@@ 5.5 Beyond English-Centric MT: The Finnish Use Case
To explore the multilingual potential of our expanded dataset via pivoting (Section 3.4) and move beyond English-centric translation, we train MT models for two additional language pairs: Finnish–Somali and Finnish-Ukrainian. The data consists of 1 876 672 sentences for Finnish-Ukranian and 1 882 712 for Finnish-Somali.
These languages were selected due to their prominence in Finland’s linguistic landscape, where Ukrainian and Somali are among the most widely spoken foreign languages, accounting for approximately 0.7% and 0.5% of the population, respectively Official Statistics of Finland (2024) (OSF). Developing high-quality models for these pairs is therefore both practical and relevant.
@@
NLLB-ft 42.21 42.31 47.61 51.62
$\Delta$ +1.88* +2.65* +2.02* +3.84*
Table 5: ChrF scores for Finnish-Somali and Finnish-Ukrainian translation.
Usefulness of Pivoted Data
@@ 6 SynOPUS: a New Synthetic Parallel Corpus Repository
The increasing adoption of LLMs in generating synthetic data underscores the growing need to systematically organize synthetic datasets. Although it is well known that many parallel datasets already contain MT content Thompson et al. (2024), when synthetic data is intentionally produced, especially when involving significant financial or computational resources, proper archiving are paramount for promoting reuse, ensuring transparency, and maximizing resource utility. Therefore, with the release of our dataset, we introduce SynOPUS, [^17] https://opus.nlpl.eu/synthetic/a new repository for parallel synthetic datasets, i.e., data that has been (partially) generated by translating text into other languages using MT systems or LLMs. We invite the community to contribute with their own datasets.
@@ 7 Conclusions
In this work, we thoroughly studied the quality and usefulness of LLM-generated synthetic data for low-resource MT. We presented a new synthetic corpus at document-level by forward translating Europarl, a parliamentary corpus, with GPT-4o. Then, we evaluated the resulting dataset both quantitatively and through human evaluation. Furthermore, we investigated the usefulness of this dataset for low-resource MT by: (i) identifying the most effective strategy for training, (ii) comparing our dataset with the public HPLT dataset, (iii) extending our analysis beyond English-centric MT by generating a multi-way parallel corpus via pivoting through alignments to English, and (iiii) studying the effect of varying training data size.
@@
Data Contamination of the Test Set
Due to the the closed-source nature of GPT-4o, there is a risk of data contamination, since the model may have already seen our test set (FLORES+) during pretraining. Recent studies Mansurov et al. (2025) have shown that distilled data may inherit intrinsic biases from the teacher model and this may have an impact on benchmark results. We note, however, that there is a strong domain mismatch between our source data (formal Europarl proceedings) and the FLORES+ benchmark (general encyclopedic text), which reduces the likelihood of simple memorization affecting results. While this risk is inherent to most widely used test sets and cannot be fully controlled, we acknowledge it here for transparency.
Data Contamination of the Source
GPT-4o pretraining data also likely includes the Europarl corpus. This means our experiments could be affected by data contamination, in the sense that the model may have had indirect prior exposure to the underlying content and domain, even if not in our low-resource target languages. Because our setup requires cross-lingual generation into languages that are not present in Europarl, direct memorization is unlikely. Still, there remains a risk of overestimating translation quality, which should be kept in mind when interpreting our results.
@@
LLMs are known to generate hallucinated content, outputs that are fluent and well-formed but factually incorrect or irrelevant Vazquez et al. (2025). This phenomenon is a risk in iself, as it can introduce noise and propagate misinformation in downstream MT models. In our case, we observed that in some cases, the model disregards the input and instead generates a response similar to, “You have been trained on data up until October 2023” in the target language. This issue is most prevalent in Georgian, with around 9,000 cases, and Ukrainian, with approximately 4,000 cases. For these languages, we removed each line containing the string “2023”. While such hallucinations appear to be an intrinsic limitation of current LLMs, they highlight the need for careful post-processing and validation when using synthetic data.
@@
Cost-Benefit Trade-off
Our empirical results demonstrate that augmenting training data with high-quality LLM-generated translations improve translation performance for low-resource languages, outperforming existing baselines. This benefit is valuable in contexts where existing parallel training data is scarce or even unavailable. However, the cost of generating such translations with LLMs is significant both in terms of compute and financial expense, and may therefore be unreasonable for many groups. For example, generating the synthetic data used in our work costed approximately \$5 000. While the improvements in translation quality may justify the use of LLMs for data generation, future work should explore more cost-efficient methods for synthetic data generation. Promising directions could include e.g., distilling larger models into smaller ones and selective data augmentation to reduce the volume of unnecessary synthetic data while preserving improvements in performance.
@@
Acknowledgments
First and foremost, we would like to sincerely thank our annotators for their valuable contributions. Ander González Docasal, Harritxu Gete Ugarte, and Nerea Mandiola Solozabal for the Basque annotation; Helga Hilmisdóttir and Reynir Eggertsson for Icelandic; Artur Voit-Antal and Yana Matyash for Ukrainian; Biljana Stojanovska for Macedonian; and Elene Kavteladze for Georgian.
We acknowledge OpenAI for their generous provision of \$5,000 in API credits, which were instrumental in generating the translations that comprise the datasets used in this work.
This project has received funding from the European Union’s Horizon Europe programme (GA No 101070350) and from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee (GA No 10052546). This work was also supported by the Digital Europe Programme under grant agreement No 101195233. and by the GreenNLP project, which is funded by the Research Council of Finland. Tiancheng Hu is supported by Gates Cambridge Trust (grant OPP1144 from the Bill & Melinda Gates Foundation).
@@ Appendix A Pilot Evaluation Results
Table 6 reports the ChrF scores of GPT-4o and EMMA on low-resource MT for a 100 random sample of the FLORES+ development set. We compare performance at the sentence (Sent.) and chunk (Chunk) level. For EMMA, both zero-shot (Sent./0) and three-shot (Sent./3) settings are reported. The best OPUS-MT model is included as a reference. There is no OPUS-MT model for Aragonese and Aranese.
We do not evaluate our synthetic baseline models on these sample sentences, as the development set was used during training. Evaluating on this data risks overfitting and leads to optimistically biased performance estimates, which may not reflect true generalization.
[…]
Table 6: ChrF scores of the evaluation of GPT-4o and EMMA on low-resource MT. Highlighted rows correspond to the final set of selected languages for our study.
@@ Appendix B Human Evaluation Annotation Guidelines
In Figure 3, we provide an exact copy of the annotation guidelines given to the annotators.
@@ Appendix C Training Regimes
  • Synthetic: We employ a shared 32k SentencePiece Kudo and Richardson (2018) vocabulary trained on the synthetic corpus; other settings follow the original Transformer-base recipe. Mini-batch fitting is enabled to optimize memory usage. Validation every 2500 updates checks perplexity. Early-stopping is employed on the development set, with a patience of 10.
@@
  • NLLB-200-distilled-1.3B-ft: We fine-tune the NLLB-distilled-1.3B model with DeepSpeed on four V100 GPUs in FP16 mixed precision. Training uses a per-GPU batch size of 32 sentences, a maximum sequence length of 128 tokens, the Adam optimiser with a 1 $\times 10^{-4}$ learning rate, and runs for up to four epochs. DeepSpeed ZeRO-1 is used for basic tensor sharding; everything else is left on-GPU. Early-stopping is employed on the development set, with a patience of 5.
@@ Appendix D OPUS-MT Models selected for fine-tuning
We select the best available OPUS-MT model based on the OPUS-MT Dashboard Tiedemann and De Gibert (2023), by looking at the BLEU score on the FLORES+ dataset.
@@ Appendix E COMET scores for MT training
We report the COMET scores for all our experiments to provide a more comprehensive evaluation. Table 7 shows the COMET scores equivalent to Table 3 for our fine-tuning experiments. Table 8 shows the COMET scores equivalent to Table 4 for our comparison with HPLT. Finally, Table 9 shows the COMET scores equivalent to Table 5 for our study on Finnish-centric translation.
Model Language Pair #\# Params
@@
$\Delta$ +29.06 +31.31 +39.67 +20.85 +39.59 +32.88 +29.94
GPT-4o 86.65 80.11 86.70 85.76 90.04 80.67 91.13
Table 7: COMET scores for our fine-tuning experiments.
Training Data Language Pair
en-eu en-is en-mk
Synthetic 81.51 80.16 82.24
HPLT 82.47 78.09 85.38
$\Delta$ +0.96 -2.07 + 3.14
HPLT 82.47 78.09 85.38
HPLT + Synthetic 84.53 82.82 86.96
$\Delta$ +2.06 +4.73 +1.58
Table 8: COMET scores for the comparison of our data with HPLT.
Model Language Pair
fi-so so-fi fi-uk uk-fi
Synthetic 75.09 66.55 76.89 77.36
OPUS-MT 54.06 34.33 89.07 88.01
OPUS-MT-ft 76.72 68.32 87.77 87.10
$\Delta$ +22.66 +33.99 -1.30 -0.91
NLLB 77.35 76.37 85.17 84.40
NLLB-ft 78.60 78.62 87.09 87.04
$\Delta$ +1.25 +2.25 +1.92 +2.64
Table 9: COMET scores for Finnish-Somali and Finnish-Ukrainian translation.
Appendix F HPLT Data Sizes
We report the total amount of sentences of the HPLT v2 dataset in Table 10 for the overlapping language pairs with our selected languages.